OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject and value of x that is not in the domain of the original logarithmic expression. Give the exact answer. log4(x+8)+log4(x-4)=3 (the 4 after the logs are subscript)

OpenStudy (anonymous):

What do you know about logs and what they *can't* be used to calculate?

OpenStudy (anonymous):

I'm confused about the part about rejecting values of x that are not in the domain.

OpenStudy (anonymous):

How do I know which value of x to use? When I take my x values and plug them back into the equation, more than one x value fits and makes everything come out correctly but I am unsure about which one to put down for answer.

OpenStudy (anonymous):

I am using MyMathLab so it's very sensitive, and I have to put in the exact correct answer

OpenStudy (anonymous):

Can you take a log of a negative number?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Good. And what about 0?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Perfect. So you know, then, that each of your logs have to be > 0. That's the limitation of your domain.

OpenStudy (anonymous):

I put in x=8 on my MathLab and got it right! You made it so easy for me, thank you!

OpenStudy (anonymous):

Yes, it is 8, but I want to make sure you know why.

OpenStudy (anonymous):

x^2-4x+8x-32=4^3 x^2+4x-32=64 -64 -64 x^2+4x-96=0 x-8=0 x+12=0 x=8 x=-12

OpenStudy (anonymous):

Oh, good! Nice job.

OpenStudy (anonymous):

(8)^2-4(8)-32=0 64-32-32=0 0=0

OpenStudy (anonymous):

Thanks. I know how to get there, just not which X to use